Heating a 6.568 g sample of an ore containing a metal sulfide, in the presence of excess oxygen, produces 1.466 L of dry SO, measured at 46.8 °C and 753.8 Torr. Calculate the percent by mass of sulfur in the ore. Assume that all sulfur in the sample was converted to SO. S(s) + O(g) → SO(g)
Answer #1

Answer:-

This question is solved by using simple concept of stoichiometry and mass percentage.

The answer is given in the image,

Answer: mairs of sample= 6.5689 V-104662 T=46.8°C =46.8+273=319.81 P =753.8 torr P=753.8 50-992 adm. 760 PV=ART pv 0.992x 1.4